OpenStudy (anonymous):

HELP MEDALS N FAN!! 1.)find the explicit rule for the nth term of the sequence 4, -4, 4, -4, ... an= 4 1^(n-1) an= 4 1^(n+1) an= 4 (-1)^n an= 4 (-1)^(n-1) 2.) find the explicit rule for the nth term of a geometric sequence where the second and fifth terms are 40 and 5000, respectively an= 8 5^(n-2) an= 8 5^(n+3) an= 8 5^(n-1) an= 8 5^(n+1)

OpenStudy (anonymous):

@amoodarya @Astrophysics

OpenStudy (amoodarya):

4 *( ?) =-4 -4 *(?)= +4

OpenStudy (amoodarya):

\[4,-4,4,-4,...\\a_n=4*(-1)^{n-1}\]

OpenStudy (welshfella):

yep thats a geometric series

OpenStudy (anonymous):

#2??

OpenStudy (anonymous):

@surjithayer

OpenStudy (welshfella):

first thin find the value of the common ratio r fifth term = ar^4 and second term = ar ar^4 / ar = r^3 so r^3 = 5000/40 find r

OpenStudy (welshfella):

can you do that r^3 = 125 r = ?

OpenStudy (anonymous):

its like 10.7

OpenStudy (anonymous):

an= 8 5^(n-1) is this one right

OpenStudy (welshfella):

no 5*5*5 = 125

OpenStudy (welshfella):

r = 5

OpenStudy (anonymous):

oh duh

OpenStudy (anonymous):

is it 5^(n-1)

OpenStudy (welshfella):

now you need to find the first term which is 40/5

OpenStudy (welshfella):

first term * 5^(n-1)

OpenStudy (anonymous):

so i was right?

OpenStudy (welshfella):

yes 8* 5^(n-1)

OpenStudy (anonymous):

thank u @welshfella !!!!

OpenStudy (welshfella):

yw